# Photovoltaic power forecasting using quantum machine learning

### v2

@@ II.5 Training and results
where $N$ is number of predictions, $\vec{x}=\left(x_{1},x_{2},\dots,x_{N}\right)$ is the predicted PV power, and $\vec{y}=\left(y_{1},y_{2},\dots,y_{N}\right)$ is the actual PV power value.
To test the models, in addition to the MSE loss metric, we also used the mean absolute error (MAE), root mean squared error (RMSE) and variance account factor (VAF).
All the machine learning simulations for this study were conducted on CPUs, on the QMware cloud qmw_qmw_2022 ; kordzanganeh2023benchmarking device. The classical part of our modeling was structured using the PyTorch library PyTorch , while the quantum part was implemented using the PennyLane framework. Notably, PennyLane provides an assortment of qubit devices. For our requirements, we selected the lightning.qubit device which is a custom backend for simulating quantum state-vector evolution. To compute the gradients of the loss function relative to each parameter: for the classical components of our hybrid models, we employed the widely-recognized backpropagation algorithm backprop ; and for the quantum part, we used the adjoint method as highlighted in Refs. Luo2020yaojlextensible ; efficient_calculation .
@@
$8-128$ 17
number of qubits $2-10$ 8
number of variational layers $1-10$ 7
@@
initial learning rate $1-1000\times 10^{-4}$ $3\times 10^{-2}$
@@
$8-128$ 3
initial learning rate $1-1000\times 10^{-4}$ $1\times 10^{-2}$
@@
$3-25$ 20
dropout range $0-99$ 0.239
number of qubits $2-5$ 4
number of variational layers $1-5$ 1
number of quantum layers $1-5$ 3
initial learning rate $1-100\times 10^{-3}$ $0.52\times 10^{-2}$
@@
$3-25$ 21
dropout range $0-99$ 0.158
initial learning rate $1-100\times 10^{-3}$ $0.5\times 10^{-2}$
Table 1: The table summarizes the hyperparameter optimization process. It shows which hyperparameters are being optimized, the range over which they could be varied, and the best values found.
@@
Table 2: Summary of the results for the proposed models. In a direct comparison, the HQNN outperforms the MLP in both training and testing losses, as evidenced across three critical metrics: RMSE, MAE, and MSE. Notably, the HQNN’s MSE is $41\%$ lower than that of the MLP. This demonstrates that the HQNN has more reliable forecasts, despite the fact that it has $1.8$ times fewer parameters. On the test dataset, the HQLSTM is $40\%$ better compared to the LTSM on the MAE and MSE metrics. It is also $21\%$ better on the RMSE metric, even though it has less than half as many weights. Overall, the HQLSTM emerges as the most precise model on all metrics. Specifically, it is $52\%$ more precise than the HQNN while having two times as few trainable parameters.
II.5.1 HQNN & MLP and HQLSTM & LSTM
Both the HQNN and its classical analog, the MLP, were trained for $20$ epochs. In contrast, the HQLSTM and its classical counterpart, the LSTM, were trained for over $50$ epochs. The Adam optimiser kingma2017adam from the PyTorch framework was used to update the parameters of the models in order to minimize their loss functions. The comprehensive training process, accompanied by the results, is shown in Fig. 4 and in Table 2.
In this study, we employed cross-validation as a fundamental technique to assess the performance of our models across distinct testing subsets. The application of cross-validation is pivotal to safeguard against potential data leakage from the training dataset into the testing dataset. To achieve this, a rigorous approach was adopted wherein a 24-hour time window, from each side of the subsets, was systematically excluded from the dataset.
@@ II.5.1 HQNN & MLP and HQLSTM & LSTM
In a head-to-head comparison between the HQNN and the MLP, the former exhibits superior performance regarding training and testing losses across three key metrics: RMSE, MAE, and MSE. Specifically, the HQNN surpasses the MLP’s power prediction accuracy by $41\%$ as estimated by the MSE loss function, and by $26\%$ on the MAE loss, all while boasting $1.8$ times fewer parameters (2266& 3987226639872266\ \&\ 39872266 & 3987).
When comparing the HQLSTM and LSTM models, the former outperforms the latter in both training and testing losses across all three of the aforementioned metrics. Remarkably, the HQLSTM has a $40\%$ better predictive ability than the LSTM as assessed by the MSE and MAE metrics. It achieves this with less than half the number of parameters (1109& 2857110928571109\ \&\ 28571109 & 2857). Moreover, HQLSTMs are more resistant to overfitting, while classical LSTMs suffer from it.
In a broader comparison encompassing all four models, the HQLSTM emerges as the most precise model on all the metrics. Notably, it is $52\%$ more precise than the HQNN, while having half as many trainable parameters.
@@
Moreover, we searched for external articles that refer to this dataset and found only one article that solves the $1$ hour ahead PV power prediction using neural networks kaloop2021novel . In comparison, our HQNN model is better than the model from this external article according to the VAF metric by $40\%$ (91& 65916591\,\&\,6591 & 65).
@@
Further, to confirm that hybrid models train better including on a smaller dataset, an additional experiment was conducted where the volume of training data was intentionally reduced. The results are shown in Figure 4. The hybrid models performed better with less data, having lower losses and better prediction capabilities than the classical models.
II.5.2 HQSeq2Seq & Seq2Seq
@@
Although the models are capable of being trained on sequences of arbitrary lengths, we chose to use sequences with a fixed $96$ hour length for simplicity. In this case, the encoder gets $96$ hours of all available features, while the decoder is asked to extrapolate only the “Power” feature of the data $96$ hours ahead. Training for $15$ epochs with the Adam optimizer (learning rate $0.001$) proved to be enough for the models to converge (Fig. 4 (f-g)).
As an example of inference, we pass the time series of the length different from $96$ into both models and prompt them to give us a forecast for $137$ hours ahead (Fig. 4 (h)). We can conclude that both models transition from the fixed sequence length to an arbitrary one quite well. It may even be possible to improve these results by introducing variable-length sequences into the training stage.
We also investigated the dependence of the test loss on the size of the training dataset for Seq2Seq and HQSeq2Seq, as shown in Fig. 4(e). As one can see, the test RMSE loss of the hybrid model is less for any size of training data, showing that the hybrid model has an advantage over the classical model, including on a substantially trimmed dataset.
III Discussion
@@
The first approach is the HQNN, a combination of classical fully connected layers and a VVRQ layer, which is a quantum analog of such layers. We compared this hybrid model with its classical counterpart, the MLP, and demonstrated that, even though the HQNN has $1.8$ times fewer variational parameters, its predictive ability is $41\%$ better, as estimated by the MSE metric.
The second approach is the HQLSTM, a hybrid quantum analogue of the classical LSTM. Here, a QDI layer is inserted into each gate of the LSTM cell. This approach provides a $40\%$ improvement in prediction using the MAE and MSE metrics compared to its classical counterpart.
The HQLSTM was even an improvement on the HQNN despite having half as many weights.
@@
We also compared our models to a paper that solved the same problem using the same dataset. We found that our best HQLSTM is $40\%$ more accurate in predicting power using the VAF metric.
@@
To fully unlock the potential of hybrid quantum approaches in time series prediction problems, further research and testing of models on other datasets is needed. More efficient optimization and training of VQCs, combined with higher fidelity and larger-scale quantum hardware, could lead to significantly greater performance improvements.
@@
While this work was done on a public dataset with an emphasis on hybrid quantum models for better forecasting performance, the quality and source of data plays a crucial role in overall effectiveness in the real world, especially considering weather data. Accurate weather forecast is a crucial input into any high performing and useful PV prediction given its dynamism and influence on PV output. An interesting area of research is cloud prediction using satellite and weather data for geolocations directly impacting solar irradiance and therefore PV output. The added complexity could enhance the need for hybrid quantum models to increase computational efficiency and higher quality forecasts.
To summarize, our developments provide three hybrid quantum approaches for time series problems that demonstrate the possibility of combining classical and quantum methods. Our proposed models show improved performance compared to classical models with similar architectures when using fewer variational parameters. We believe that these results pave the way for further research in developing hybrid models that leverage the strengths of both classical and quantum computing.
@@
Data availability All datasets used in this study are publicly available and are cited within the Dataset subsection of the paper. The machine learning models employed are fully described in the Results section, enabling independent reproduction of the results.

### v3

@@ III Results and discussion
where $N$ is the number of predictions, $\vec{x}=\left(x_{1},x_{2},\dots,x_{N}\right)$ is the predicted PV power, and $\vec{y}=\left(y_{1},y_{2},\dots,y_{N}\right)$ is the actual PV power value.
To test the models, in addition to the MSE loss metric, we also used the mean absolute error (MAE), root mean squared error (RMSE), variance accounted for (VAF), mean absolute percentage error (MAPE), and coefficient of determination ($R^{2}$).
The application of the squaring operation in the calculation of the RMSE results in the amplification of outlier influence, assigning significant weight to rare but substantial errors. Such errors can adversely affect power grid balancing processes and reserve capacity sizing. In contrast, the MAE and the MSE characterize the overall pointwise accuracy of a model and often demonstrate greater relevance for energy tasks. The VAF and the $R^{2}$ quantitatively assess the model’s explanatory power by measuring the explained variance (higher values are preferable; $\text{VAF}=100\%$ or $R^{2}=1$ corresponds to a perfect fit). The MAPE provides a dimensionless measure of error in percentage terms, facilitating its operational interpretation; however, this metric may demonstrate instability when true values approach zero.
All the machine learning simulations for this study were conducted on CPUs using the QMware cloud qmw_qmw_2022 ; kordzanganeh2023benchmarking device. The classical part of our modeling was structured using the PyTorch library PyTorch , while the quantum part was implemented using the PennyLane framework. For our requirements, we selected the lightning.qubit device which is a custom backend for simulating quantum state-vector evolution. To compute the gradients of the loss function relative to each parameter: for the classical components of our hybrid models, we employed the widely-recognized backpropagation algorithm backprop ; and for the quantum part, we used the adjoint method as highlighted in Refs. Luo2020yaojlextensible ; efficient_calculation .
@@
88128128 17
number of qubits $2\text{--}10$ 8
number of variational layers $1\text{--}10$ 7
@@
initial learning rate $1\text{--}1000\times 10^{-4}$ $3\times 10^{-2}$
@@
$8\text{--}128$ 3
initial learning rate $\text{--}1000\times 10^{-4} 1×1$$\times 10^{-2}$
@@
$3\text{--}25$ 20
dropout range $0\text{--}99$ 0.239
number of qubits $2\text{--}5$ 4
number of variational layers $1\text{--}5$ 1
number of quantum layers $1\text{--}5$ 3
initial learning rate $1\text{--}100\times 10^{-3}$ $0.52\times 10^{-2}$
@@
$3\text{--}25$ 21
dropout range $0\text{--}99$ 0.158
initial learning rate $1\text{--}100\times 10^{-3}$ $0.5\times 10^{-2}$
Table 1: Summary of the hyperparameter optimization process for the HQNN, MLP, HQLSTM and LSTM models. It shows which hyperparameters are being optimized, the range over which they could be varied, and the best values found.
@@
test, MAPE $\times 10^{6}$ $33.5$ $32.2$ $26.6$ $40.8$
test, VAF $89.57$ $89.71$ $92.01$ $86.64$
test, $R^{2}$ $0.8703$ $0.8767$ $0.9190$ $0.8338$
Table 2: Summary of the results obtained by the HQNN, MLP, HQLSTM and LSTM models on the Mediterranean PV dataset malvoni2016data .
III.1 HQNN versus MLP, and HQLSTM versus LSTM
Both the HQNN and its classical analogue, the MLP, were trained for $20$ epochs. In contrast, the HQLSTM and its classical counterpart, the LSTM, were trained for over $50$ epochs. The Adam optimizer kingma2017adam from the PyTorch framework was used to update the parameters of the models in order to minimize their loss functions. The comprehensive training process, accompanied by the results, is shown in Fig. 4 and in Table 2.
In this study, we employed cross-validation as a fundamental technique to assess the performance of our models across distinct testing subsets. The application of cross-validation is pivotal to safeguard against potential data leakage from the training dataset into the testing dataset. To achieve this, a rigorous approach was adopted in which a 24-hour time window on each side of the subsets was systematically excluded from the dataset.
@@ III.1 HQNN versus MLP, and HQLSTM versus LSTM
To verify the quality of the quantum layers assembled by Optuna, a theoretical analysis of quantum circuits was performed. Evaluation of ZX-calculus metrics, Fisher information, and Fourier series metrics showed that the resulting models possess high expressibility, trainability, and resource efficiency. The analysis is described in detail in Section IV.
After hyperparameter optimization, the total number of parameters for the models was $1109$, $2266$, $2857$, and $3987$ for the HQLSTM, HQNN, LSTM, and MLP models, respectively. The HQLSTM has the fewest parameters (approximately $39\%$ of its classical counterpart, the LSTM) and the HQNN has the second fewest parameters (approximately $57\%$ of its classical counterpart, the MLP).
In a head-to-head comparison between the HQNN and the MLP, the former exhibits superior performance on 5 out of the 8 metrics. The largest training improvement was the $59\%$ reduction in the MSE, and the largest testing improvement was the $41\%$ MSE reduction (see Table 2).
When comparing the HQLSTM and LSTM models, the HQLSTM was better on 7 out of the 8 metrics. Most significantly, both the test MAE and test MSE improved by about $40\%$. Moreover, the HQLSTM was more resistant to overfitting than the classical LSTM.
Regarding the two hybrid quantum models, the HQLSTM outperformed the HQNN in every metric. Its test MSE and test MAE scores were $52\%$ and $48\%$ better, respectively, while having half the number of trainable parameters. In a broader comparison encompassing all four models, the HQLSTM emerged as the best model on 7 out of the 8 metrics. This is especially notable since the HQLSTM has the fewest trainable parameters of the four models.
Since the HQLSTM model demonstrates competitive performance, we further evaluated it against the established classical ensemble methods, XGBoost and LightGBM, to contextualize its effectiveness. The HQLSTM has orders of magnitude fewer trainable parameters than these gradient-boosting tree algorithms.
As shown in Table 3, the HQLSTM model achieves the best values for 2 of the 7 metrics, RMSE and MAPE. There are negligible differences in the VAF and $R^{2}$ values, and the boosting models lead in MAE and MSE.
This is due to known differences in each metric’s sensitivity: RMSE penalizes occasional large deviations more heavily, which are particularly disruptive for grid-balancing and reserve-capacity planning. Conversely, MAE and MSE capture overall point accuracy and may be preferable for revenue-settlement tasks. These findings highlight the importance of metric selection in model evaluation, with HQLSTM offering a compelling alternative for applications prioritizing error distribution robustness over absolute precision.
@@
It is also worth noting that the high value of the MAPE in all models is associated with a large number of zero values of the target variable during nighttime. Importantly, the HQLSTM reaches ensemble-level accuracy and does so with orders of magnitude fewer trainable parameters. Moreover, when we searched for external articles that refer to this dataset, we found only one article that solved the $1$ hour ahead PV power prediction using neural networks kaloop2021novel , and our HQLSTM model was better than the model from this article by $40\%$ ($92$ versus $65$) according to the VAF metric.
Model HQLSTM XGBoost LightGBM Gain
test, MAE $0.0266$ $0.0260$ $0.0251$ $3.46\%$
test, MSE $0.0033$ $0.0030$ $0.0030$ $9.09\%$
test, RMSE $0.0539$ $0.0551$ $0.0550$ $2.00\%$
test, MAPE $\times 10^{6}$ $13.1$ $20.7$ $14.4$ $9.03\%$
test, VAF $93.08$ $93.75$ $93.77$ $0.02\%$
test, $R^{2}$ $0.9307$ $0.9373$ $0.9375$ $0.02\%$
Table 3: Test performance comparison between the proposed HQLSTM model and classical ensemble methods such as XGBoost and LightGBM. For convenient interpretation of these results, the “Gain” column shows the relative improvement from the second best to the best value of the performance metric. The data was obtained for runs without k-fold validation.
To further evaluate the generalization capability of HQLSTM beyond the original PV power prediction task, we conducted additional experiments using the Appliances Energy Prediction dataset appliances_energy_prediction_374 . This dataset involves forecasting energy consumption in a smart home environment based on multiple contextual variables such as indoor temperature and humidity, outdoor weather conditions, and time-related features.
To ensure compatibility with our modeling pipeline, the original 10 minute resolution data were aggregated to hourly intervals and transformed into fixed-size time windows. Specifically, the model received 3 hours of past data to predict energy usage 1 hour ahead. Eleven features with the highest correlation to the target variable were retained for training.
Both classical LSTM and HQLSTM models were retrained under the new settings with the HQLSTM having $20622$ trainable parameters and the LSTM having $89257$. A head-to-head comparison is shown in Table 4 and Fig. 5, where it can be seen that the HQLSTM excels in 3 out of the 6 performance metrics. The largest relative improvement is the $12.5\%$ reduction in the MAPE, which shows that the HQLSTM has better control over its typical and relative errors.
Model HQLSTM LSTM Gain
test, MAE $0.0513$ $0.0535$ $4.11\%$
test, MSE $0.0090$ $0.0089$ $1.11\%$
test, RMSE $0.0937$ $0.0934$ $0.32\%$
test, MAPE $38.39$ $43.88$ $12.51\%$
test, VAF $38.07$ $37.83$ $0.63\%$
test, $R^{2}$ $0.3595$ $0.3682$ $2.42\%$
Table 4: Performance comparison between HQLSTM and LSTM on the additional energy consumption dataset. As before, “Gain” is the relative improvement from the second best to the best value of the performance metric.
These results suggest that HQLSTM is more robust to individual prediction errors, despite using $4$ times fewer trainable parameters. They also confirm that the proposed HQLSTM architecture generalizes well to new forecasting tasks and retains strong performance characteristics across domains.
Figure 5: Training dynamics of HQLSTM and LSTM models on the energy prediction task. (a) MSE loss on the training subset shows both models converging similarly. (b) MAE loss on the test subset indicates that HQLSTM achieves lower test loss values overall. Specifically, the lowest loss is $0.0513$ and $0.0535$ for the HQLSTM and LSTM models, respectively.
@@
Further, to confirm that hybrid models train better, including on a smaller dataset, an additional experiment was conducted where the volume of training data was intentionally reduced. The results are shown in Figure 4(d–e). The hybrid models performed better with less data, having lower losses and better prediction capabilities than the classical models.
III.2 HQSeq2Seq versus Seq2Seq
@@
Although the models are capable of being trained on sequences of arbitrary lengths, we chose to use sequences with a fixed $96$ hours length for simplicity. In this case, the encoder gets $96$ hours of all available features, while the decoder is asked to extrapolate only the “Power” feature of the data $96$ hours ahead. The Seq2Seq model had 2705 trainable parameters and the HQSeq2Seq model had 2750.
Model training was conducted using the Adam optimizer (learning rate $0.001$) with a maximum of 15 epochs. We additionally applied early stopping based on test loss, and observed that both models typically converged within the first $13\text{--}15$ epochs (see Fig. 4(f–g)). As such, extending training further provided no meaningful performance gains, while increasing computational cost.
As an example of inference, we pass the time series of a length different from $96$ into both models and prompt them to give us a forecast for $137$ hours ahead (Fig. 4(h)). Both models generalize well from the fixed sequence length to an arbitrary one. It may even be possible to improve these results by introducing variable-length sequences into the training stage.
We also investigated the dependence of the test loss on the size of the training dataset for Seq2Seq and HQSeq2Seq, as shown in Fig. 4(e). As one can see, the test RMSE loss of the hybrid model is lower for any size of training data, showing that the hybrid model has an advantage over the classical model, including on a substantially trimmed dataset.
III.3 Discussion
@@
The first approach is the HQNN, a combination of classical fully connected layers and a VVRQ layer, which is a quantum analogue of such layers. We compared this hybrid model with its classical counterpart, the MLP, and demonstrated that, even though the HQNN has $1.8$ times fewer variational parameters, its predictive ability is $41\%$ better, as estimated by the MSE metric.
The second approach is the HQLSTM, a hybrid quantum analogue of the classical LSTM. Here, a QDI layer is inserted into each gate of the LSTM cell. This approach provides a $40\%$ improvement in prediction using the MAE and MSE metrics compared to its classical counterpart. The HQLSTM was even an improvement on the HQNN despite having half as many weights.
@@
We also compared our models with those reported in a paper that solved the same problem using the same dataset. We found that our best HQLSTM is $40\%$ more accurate in predicting power using the VAF metric.
To complement our performance evaluation, we conducted an architectural analysis of the quantum layers used in our models. As detailed in Section IV, we compared the VVRQ and QDI layers using ZX-calculus, Fisher information matrices, and Fourier spectral analysis. These methods allowed us to assess the efficiency, trainability, and expressive capacity of the circuits beyond their end-task performance. The results reveal concrete trade-offs. VVRQ is ZX-irreducible (all $56/56$ parameters retained) and shows near-uniform FIM spectra without rank saturation as depth grows, indicating depth-scalable trainability; its Fourier analysis (approximate, due to size) indicates an increasing number of non-zero coefficients with qubit count, consistent with broad expressivity. In contrast, QDI is also ZX-irreducible ($24/24$ parameters) but achieves high trainability at shallow depth: its FIM enters the over-parameterized regime (rank saturation), while its Fourier spectrum is densely populated ($151/161\approx 94\%$ non-zero coefficients), making it compact and well-suited for resource-constrained hybrids such as HQLSTM. Taken together, these measurements clarify the design space of quantum components and support informed choices between depth-scalable VVRQ and lightweight yet expressive QDI in future hybrid QML architectures.
@@
To fully unlock the potential of hybrid quantum approaches in time series prediction problems, further research and testing of models on other datasets are needed. More efficient optimization and training of VQCs, combined with higher fidelity and larger-scale quantum hardware, could lead to significantly greater performance improvements. While the study focuses on performance characteristics of hybrid quantum-classical models, we acknowledge that their training and tuning can be more complex than those of purely classical counterparts. This is due to the need for domain knowledge in both machine learning and quantum computing, as well as simulator-related constraints. Future work may explore more accessible training workflows and improved tool-kits for hybrid model development.
@@
While this work was done on a public dataset with an emphasis on hybrid quantum models for better forecasting performance, the quality and source of data play a crucial role in overall effectiveness in the real world, especially considering weather data. Accurate weather forecasts are a crucial input to any high performing and useful PV prediction given its dynamism and influence on PV output. An interesting area of research is cloud prediction using satellite and weather data for geolocations directly impacting solar irradiance and therefore PV output. The added complexity could enhance the need for hybrid quantum models to increase computational efficiency and to provide higher quality forecasts.
III.4 Conclusion
In this work, we addressed the important problem of photovoltaic time-series forecasting. This is a crucial task for successfully operating solar power stations and managing their integration with the electricity grid. To this end, we proposed three hybrid quantum–classical time-series models, namely HQNN, HQLSTM and HQSeq2Seq, and evaluated them on a dataset from a Mediterranean power plant. We used five-fold cross-validation to carefully analyze their performance relative to their most comparable classical counterparts. We made additional comparisons with XGBoost and LightGBM, performed tests where we deliberately reduced the size of the training data, and assessed whether the HQLSTM model could generalize to an energy-consumption prediction task. Our contributions can be divided into three parts. First, we integrated quantum circuits (VVRQ and QDI) into standard MLP, LSTM, and Seq2Seq backbones. Second, we developed an HQSeq2Seq model that can forecast user-defined horizons without requiring future meteorological inputs. Third, we included a circuit-level analysis based on ZX-calculus, Fisher information, and Fourier metrics.
[…]
Future work could include validating the models across additional locations and with higher-frequency settings, developing more efficient and user-accessible approaches to the training of variational quantum circuits, and evaluating performance on actual (rather than simulated) quantum devices.
@@ III.4 Conclusion
Data availability
All datasets are publicly available and are cited within the Dataset subsection of the paper. The machine learning models employed are fully described in the Results section, enabling independent reproduction of the results.